MODS1 Commissioning Run 6: UTC 2011 March 21

Observer(s): R. Pogge (OSU), P. Garnavich (ND), and M. Pedani (LBTO)
Telescope Operator: S. Allanson (LBTO)
AO Support: as-needed from Tucson
Telescope Support: M. Midkiff (LBTO)
Instrument Support: D. Officer (LBTO)


Worked the first two hours of the night before high winds shut us down.

Success testing the newest iteration of the script. Got in our tests before wind chased us away from the Stone F field. Got one long-slit spectral acquisition done before the wind went over limits and shut us down.


Paul Martini examined the Stone F MOS mask data from last night and sent us an updated version of (v0.3.1). Using this on last night's data shows it has cleared up most of the issues we encountered. We are ready to try it on-sky tonight.


All times are UTC unless otherwise noted.

0005 - Thick high clouds have covered a fair fraction of the sky, coming in from the NW. The satellite shows a hole approaching from over Baja so we might get some clearing before the main clouds come in. The wind is our biggest concern, we've lately had gusts up to 23m/s, so we'll be watching that and the clouds.

0104 - Clouds are getting thinner overhead, but winds remain gusty, up to 24m/s in the past 10 minutes.

0145 - Wind slowly trending downward, and the clouds have thinned a little as the main wave of clouds is starting to pass us by. A lot of junk still to the WNW, but if the wind drops enough we may be able to open.

0150 - Winds are still declining and not a gusty. Ready to open. We're going straight to Stone F after pointing and baseline collimation.

0205 - Pointing star
   IE = -70
   CA = +86

Stone F MOS Test Field

0210 - Going to the Stone F MOS Test Field

Paul Martin is connected via video skype while we test the mask

Acquisiton with acqMODS and the stone_f.acq script. Worked just fine. Since we have not yet gone to a collimation star, we took longer to converge, but not too bad.

Thru-mask and Field images from the acqMODS script.

Ran --> dX=-0.56 dY=-0.47 dPA=-0.2491

Executed offsetpointing -0.2491 -0.56 -0.47 detxy rel left

Thru-mask image. Running -s to get shift-only to take out residual shift - rotation looks good

found dX=-0.05 dY=-0.31, consistent with the residual misalignment I see by-eye (OK in X, a little high in Y).

But, even without this if these were slits, the targets would be on the center of the slit in dispersion and only a little displaced in Y (2.5 pixels).

Final image after applying that shift (offsetxy -0.05 -0.31 rel) the stars are beautifully centered to within the expected errors.

With the well-aligned image, added a rotation offset of dPA=0.25 degrees to use for further refinement later.

I think this means we have the problems from the last night fixed. The acquire and align process went as we expected and worked fine. With a few more tweaks, it should even be reasonably efficient.

0235 - A pause in the action while Paul and I review the results, Steven pointing to a star away from the wind to watch the wind speed.

0248 - Back to Stone F

0306 - getting spikes on azimuth torques from the wind, stopping observing

watching the wind

0325 - Target NGC 2315, long-slit acq test

On-target by 03:26, offset to slit center at 03:32, start of spectrum acquisition at 03:34, remembering we're doing this by-hand still

Seeing 0.9-1.1 arcsec. Still windy, but we're getting shadowed by the builidng (about 135° relative to the wind). Watching the gusts, we could still get over the limit. Moon is rising, illuminating clouds - we're shootting through cirrus

0353 - Wind over limits, exposures stopped a little after 300s and read out. Steve closing the doors.

0515 - Wind gusts now getting up to 25m/s, but clearing overhead for now, with more stuff on the way...

0800 - Sustained 20m/s winds for the past couple hours, with gusts up to nearly 30m/s and clouds blowing past rapidly.

We're calling the night early, especially as two of the observers (Peter and Marco) will be getting up early to leave the mountain before the forecast snow arrives tomorrow.

-- RichardPogge - 20 Mar 2011
Topic revision: r6 - 21 Mar 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback